# Quantum Imaging and Metrology with Undetected squeezed Photons: Noise Canceling and Noise Based Imaging

### v1

@@ IV.2 The Limit $x\rightarrow 1$
Substituting (16) into (7) and differentiating the resulted equation with respect to $\phi_{T}$ leads to,
@@
It is worth mentioning that in Eq. (17), we take $\delta=k\pi/2$. The sensitivity given by Eq. (17) is independent of $x$ and proportional to the factor $G_{1}G_{2}$. We substitute (16) and its corresponding $\phi_{G}$ in Eq.(11) to obtain SNR,
$\text{SNR}=\begin{cases}2\alpha^{2}\sin^{2}\left(\frac{\Delta-\phi_{T}}{2}\right)&k\,\,\text{is odd}\\ 2\alpha^{2}\cos^{2}\left(\frac{\Delta-\phi_{T}}{2}\right)&k\,\,\text{is even}\,.\end{cases}$ (18)
Since Eq.(18) depend on adjustable parameter $\alpha$, SNR one can improve it by increasing the mean photon numbers in input signal field of NOPO.
V Imaging Protocol
This section is devoted to how construct the image using either of quantum fluctuation or signal. In the following subsection we present a protocol to extract object information by quantum fluctuation. Also the image construction protocol using the homodyne signal is the subject of subsection V.2
V.1 Imaging Protocol Using Quantum Fluctuation
Since the quantum fluctuation (noise term) depends on the parameter $G$ (see Eq. (10)), which contains the object information, we can construct the image using the quantum fluctuation. To this end we use Eq.(9) to rewrite Eq.(10) as follows,
@@ V.1 Imaging Protocol Using Quantum Fluctuation
It is worthwhile to emphasis that $\Delta$ is experimentally adjustable parameter via precise tuning the phase mismatch of the classical fields pumping the crystals, see figure 1. Therefore the image construction processes can be achieved by the following protocol:
i) At first stage, the values of $\langle\Delta S^{2}(\Delta)\rangle$ for $\Delta=0,\pi/2,3\pi/2,\pi$ is measured.
ii) The 2nd stage is to calculate the image information by the following formulas:
@@
V.2 Imaging Protocol using Signal
@@
VI Summary and Conclusion
In the present work an imaging setup is considered. In the proposed setup a classical pump field excites a non-linear crystal posses second order susceptibility, $\chi_{1}^{(2)}\neq 0$, to produce non-degenerate squeezed photon pair (called signal and idler) which is aligned on the center of a confocal cavity to construct a non-degenerate optical parametric oscillatorer (NOPO) . The NOPO is pumped by a coherent light and stimulated by a input coherent beam prepared in the signal frequency. The output idler beam emerging from NOPO interacts with the object and then injected into a second non-linear crystal, same to the first crystal. This idler beam merges with idler beam generated in the second crystal during SPDC process, in order to achieve path identity. The image information can be extracted by a balanced homodyne detection of the signal beam emerged from NOPO and signal beam generated by the SPDC process in the second crystal. Notice that, these signal beams never interacts with the object. Along with the theoretical treatment of the setup, the quantum fluctuation is also calculated. Interestingly, we found that the image information (transition coefficient and phase) can be also extracted from quantum fluctuation. Our calculations show that SNR is enhanced as a function of the photon mean number of input signal field. The further theoretical calculation is applied for both oblique object ($x\rightarrow 0$) and transparent object($x\rightarrow 1$) cases.
For the first regime ($x\rightarrow 0$) the behavior of SNR strongly depends on the relative phase of the input signal field to the NOPO and homodyne local oscillator. In this regime, at the high and low parametric gains, SNRs are scaled by $0.5$ and $1$ respectively. The second regime occurs at high parametric gain and high object transmission ($x\rightarrow 1$). In this regime, SNR is independent of parametric gains.
Finally we presented an imaging protocol to remove classical background noises. In this protocol, the output electric signals from the balanced homodyne setup are measured after adjusting phases $\phi_{\alpha}$, $\phi_{\beta}$ and the pumps phases as described in Sec.V. In this way, the image information can be superimposed on the homodyne signal, where background noises are suppressed. The noiseless image information (such as object transmission and phase) can be extracted from the signals which are measured by the balanced homodyne setup.
In view of the aforementioned point the results of the present work provide effective theoretical treatment to optimize SNR in the QUIP. Moreover the proposed setup and the presented imaging protocol lead to QUIP without background noises.

### v2

@@ 3.2 High gain, High Transmission, $x\rightarrow 1$
Substituting (17) into (7) we get the following for the case of $\delta=k\pi/2$.
@@
In this case the sensitivity is independent of $x$ and proportional to the factor $G_{1}G_{2}$. The $\text{SNR}_{HD}$ becomes:
$\text{SNR}_{HD}=\begin{cases}2\alpha^{2}\sin^{2}\left(\frac{\Delta-\phi_{T}}{2}\right)&k\,\,\text{is odd}\\ 2\alpha^{2}\cos^{2}\left(\frac{\Delta-\phi_{T}}{2}\right)&k\,\,\text{is even}\,,\end{cases}$ (19)
which is dependent on adjustable parameter $\alpha$. Hence, $\text{SNR}_{HD}$ can be improved by increasing the mean photon numbers in input signal field of NOPO.
4 Imaging Protocols
In this section we introduce two data acquisition protocols for image construction. First we present a protocol to extract object information by quantum fluctuation (noise term).
4.1 Imaging Protocol Using Quantum Fluctuation
Since the quantum fluctuation (noise term) depends on the parameter $G$ (see expression for $\langle\Delta S^{2}\rangle$), which contains the object information, we can construct the image using the information encoded into quantum fluctuation. To this end we use Eq.(9) to rewrite $\langle\Delta S^{2}\rangle$ as follows:
@@ 4.1 Imaging Protocol Using Quantum Fluctuation
It is worthwhile to emphasis that $\Delta$ is experimentally adjustable parameter via precise tuning the phase mismatch of the classical fields pumping the crystals, see figure 1. Therefore the image construction processes can be achieved by the following steps:
i) At first step, the values of $\langle\Delta S^{2}(\Delta)\rangle$ for $\Delta=0,\pi/2,3\pi/2,\pi$ is measured.
ii) The 2nd step is to calculate the image information by the following formulas:
@@
Here we have provided a physical explanation of image formation resulting from noise measurements. In the next section introduce a protocol to extract image information from output signal of homodyne detection.
4.2 Imaging Protocol using Signal
@@
5 Summary and Conclusion
We have presented an imaging setup with undetected photons. The resource photons (signal and idler photon pairs) are generated in two non-linear crystals (posses second order susceptibility, $\chi_{1}^{(2)}\neq 0$) via SPDC process. In order to enhance the imaging frame rate and reduction of undesired influences of noise (improving SNR), each nonlinear crystal is placed within a confocal cavity, forming a non-degenerate optical parametric oscillatorer (NOPO). The first NOPO is pumped by a coherent beam and seeded with a coherent signal beam. The second NOPO is pumped by a coherent beam (with certain phase relation to the first pump beam) and seeded by signal and idler beams emerging from first NOPO to ensure path identity. The image information can be extracted by a balanced homodyne detection of the signal beams emerged from two NOPOs. Notice that, these signal beams never directly interacts with the object. Our extended theoretical treatment, valid beyond the low gain approximation, demonstrates that the SNR improves with increasing mean photon number in the input seed signal. Detailed analysis for the limits $x\rightarrow 0$ (opaque object)and $x\rightarrow 1$(transparent object) reveals that the sensitivity depends on the relative phase of the input seed signal and local oscillator in the former case, while it becomes independent of the parametric gains in the latter. We have also introduced imaging protocols that extract object information from both quantum fluctuation (noise) and the homodyne signal, effectively suppressing background noise.
These results provide a robust theoretical framework for optimizing SNR in quantum imaging with undetected photon (QUIP) and pave the way for noise-free imaging applications in sensitive bio-imaging.